OpenStudy (czarluc):

find the solution sets for ->secx+tanx=0 ->sinxtanx=0 ->sin2x+sin3x=0

OpenStudy (perl):

are these two separate equations?

OpenStudy (czarluc):

yes

OpenStudy (czarluc):

would be great if all are answered :)

OpenStudy (anonymous):

Could you edit them to show it's 2 because im confused.

OpenStudy (czarluc):

actually there are three questions lol

OpenStudy (solomonzelman):

\(\large\color{black}{ \sec x+\tan x=0}\) \(\large\color{black}{ \sec x=-\tan x}\) square both sides, \(\large\color{black}{ \sec^2x=\tan^2x}\) \(\large\color{black}{ \tan^2x+1=\tan^2x}\) your opinion?

OpenStudy (czarluc):

then it would be 1=0?

OpenStudy (solomonzelman):

yes, and that means that how many solutions are there?

OpenStudy (czarluc):

1? sorry I don't know :(

OpenStudy (solomonzelman):

No solution, because 1 can't equal zero. (No?)

OpenStudy (solomonzelman):

No solution, because 1 can't equal zero. (No?)

OpenStudy (czarluc):

the solutions sets I 'm referring to are the angles like in 0<=x<=360

OpenStudy (czarluc):

how do you solve this in the forms of trig. equations?

OpenStudy (solomonzelman):

\(\large\color{black}{ \sin(3x)+\sin(2x)=0}\) \(\large\color{black}{ \sin(3x)=-\sin(2x)}\) \(\large\color{black}{ \sin(3x)=\sin(-2x)}\) \(\large\color{black}{ 3x=-2x}\)

OpenStudy (solomonzelman):

So x=?

OpenStudy (czarluc):

0

OpenStudy (solomonzelman):

I am not so sure about what the second question is, because it is not equal to anything. You have to right the second question correctly.

OpenStudy (solomonzelman):

And yes, x=0 for #3.

OpenStudy (solomonzelman):

Oh you did already:)

OpenStudy (czarluc):

but when i look it up on the net they say it has like 7 solutions

OpenStudy (czarluc):

hahaha thanks

OpenStudy (czarluc):

or maybe not seven just not only zero

OpenStudy (solomonzelman):

\(\large\color{black}{ \tan(x)\sin(x)=0}\) \(\large\color{black}{ \tan(x)=0}\) or \(\large\color{black}{ \sin(x)=0}\)

OpenStudy (solomonzelman):

There can be many solutions. Are you given any intervals?

OpenStudy (czarluc):

uhm its 0<=x<=360

OpenStudy (solomonzelman):

Ohh, so, [0,360]

OpenStudy (solomonzelman):

When is sin(x)=0 ? when x is what, can you tell me?

OpenStudy (solomonzelman):

There are 2 x-solutions. 180º and 360º

OpenStudy (solomonzelman):

those questions marks are supposed to be degree signs, nothing more...

OpenStudy (czarluc):

sorry for the delay the internet was cut off

OpenStudy (solomonzelman):

happens, it's fine:)

OpenStudy (czarluc):

would you include 0?

OpenStudy (solomonzelman):

yes, since it is a closed interval, the zero is included.

OpenStudy (czarluc):

oh so there are three answers?

OpenStudy (solomonzelman):

yes

OpenStudy (czarluc):

oh ok got it! thanks :)

OpenStudy (solomonzelman):

Enjoy...